OpenStudy (undercover):

8x+y=-16 -3x+y=-5 solve the systems by elmination Name the solution as an ordered pair

OpenStudy (anonymous):

I dont know if this is right, but if you are solving for Y it would look like this 8x+y=16 -8x+y=16-8x so, Y+16-8x?